OpenStudy (anonymous):

which of these inequalities is equivalent to 2x - 8y less than or equal to -3

HanAkoSolo (jamierox4ev3r):

\(2x-8y\le -3\) right?

OpenStudy (anonymous):

yes

HanAkoSolo (jamierox4ev3r):

I'm not sure what your other inequalities are, but in most cases, you can treat the \(\le\) as an equal sign and solve for whatever variable you want. Just remember, when you divide by something negative, then the sign changes

hartnn (hartnn):

prolly one of the options says -2x+8y >= 3

OpenStudy (anonymous):

they all v 1/4x - 3or 3/8

hartnn (hartnn):

ok, so they isolated y 2x-8y <=-3 start by subtracting 2x from both sides

OpenStudy (anonymous):

wha was you4 answer?

HanAkoSolo (jamierox4ev3r):

\(2x-8y\le -3\) -2x -2x ------------- \(-8y\le -2x-3\) \-8 \-8 \-8 --------------- \(y\ge ??\) ^^It's just like solving a regular equation, but note that the sign changes since you are dividing by -8